How to find percentage decrease in the weight of the body when take to height 32 km above the surface of earthif radius of earth is 6400 km...Ans is 1 % how?

Grade:11

2 Answers

Vishal
7 years ago
g(h)= g(1-2h/r) { g(h)- gravity at height h % decrease= 1 Kitna likhu bhai? Samjh ja ab I am anonymous trying to help but these people are not allowing me to do so.
Sonia Paruthi
6 years ago
Here d= 32 km R= 6400 kmweight of body at depth d is mg` =mg(1-d/R)% decrease in weight =( mg- mg`)/mg × 100= 32 / 6400×100 = 0.5%.